Problem 5: Consider the initial value / Dirichlet problem ut(t, x) – 2uzz(t, x) = e, (t, x) € (0, +00), u(0,x) = 1, u(t,0) = please only use substaction method or method of recflection
Answer #1


u (t, x) U(0,0)=1 + 2 Uxa (t, x) = e-t ult, o) e-t (tn) € (0,700) equation Ig Iz ho moyenous boundary This heat uttro) and uf

Please do like